Project

General

Profile

PC-2015-11-19-diagnostics » History » Version 6

Rogers, Chris, 19 November 2015 16:13

1 1 Rogers, Chris
h1. Diagnostics
2
3
There will be an analysis/diagnostics meeting Thursday 19th November, 14:00.
4
5
We will be in one of the usual meeting rooms in the Atlas building, using phone:
6
7
http://mice.iit.edu/phonebridge.html
8
9 5 Rogers, Chris
* Tracker resolution and efficiency - Melissa/Chris Hunt/John Nugent
10 1 Rogers, Chris
* Tracker-to-tracker alignment - Melissa/Chris Hunt/John Nugent
11
** Tracker Survey
12
* Tracker alignment to field - Chris Rogers/Sophie
13
* Cut-based beam selection vs 2015/01 data - Tanaz/Scott Wilbur
14
* Ckov momentum scan - Ao, Durga
15
* Full PID vs 2015/01 data - Celeste
16 3 Drielsma, Fran├žois
* PID detector alignment - Francois
17 1 Rogers, Chris
* Detector Efficiencies - Jan/Francois
18
* Paper: High precision measurement of emittance - Chris Hunt/Victoria
19 6 Rogers, Chris
20
h2. Notes
21
22
Celeste Pidcott
23
Francois Drielsma
24
Chris Rogers
25
Tanaz Mohayai
26
Ryan Bayes
27
John Nugent
28
Scott Wilbur
29
30
Millipede offsets etc fixed
31
- please show residuals plots
32
- confirming internal alignment
33
- getting ready for tracker to tracker alignment for next time
34
35
Francois, tracker sign flip (Drielsma)
36
- xflip - reading out fibres the wrong way for all the planes
37
- v and w were swapped
38
- wrong ordering of the readout is at reconstruction level
39
- plane swap comes from the geometry
40
- still need to verify fix in MC
41
42
Tracker alignment to field (VB)
43
- VB promises for something today or tomorrow(ish)
44
45
Cut-based beam selection (Scott)
46
- Scott has some working code that will do a framework for cut-based beam selection
47
- Do we want to run it inside the MAUS loop or just get it running?
48
- Figure out mapping from D2 current to timing windows
49
- Notes that D2 currents are not necessarily related in the expected way to TOF012; will turn into powerpoint for next time
50
51
Cut-based beam selection (Tanaz)
52
* Noticed that discrepancy between TOF peaks between summer and October runs
53
54
Likelihood PID (Celeste)
55
* Run PID on 2015 user cycle
56
* Things look like they are working well
57
* Nothing quantitative yet
58
* Things look reasonable
59
* Only been done for KL and TOF; EMR not MCing and Ckov not in framework yet
60
* Not yet looking at consistency between U/S and D/S as there are not yet enough detectors in place
61
62
Detector alignment
63
* Nothing more
64
* Could improve by using different tracker stations
65
* Need to correct for tracker pitch
66
* Looking at fitting distribution centre rather than calculating naive mean/RMS